OpenStudy (anonymous):

help! What is the value of g[f(3)] for the functions f(x) = 2x + 3 and g(x) = x + 1?

jimthompson5910 (jim_thompson5910):

what is f(3) equal to?

jimthompson5910 (jim_thompson5910):

f(x) = 2x + 3 f(3) = 2(3) + 3 f(3) = ???